# Contrast Sets for Evaluating Language-Guided Robot Policies

### v1

@@ Abstract
Robot evaluations in language-guided, real world settings are time-consuming and often sample only a small space of potential instructions across complex scenes. In this work, we introduce contrast sets for robotics as an approach to make small, but specific, perturbations to otherwise independent, identically distributed (i.i.d.) test instances. We investigate the relationship between experimenter effort to carry out an evaluation and the resulting estimated test performance as well as the insights that can be drawn from performance on perturbed instances. We use contrast sets to characterize policies at reduced experimenter effort in both a simulated manipulation task and a physical robot vision-and-language navigation task. We encourage the use of contrast set evaluations as a more informative alternative to small scale, i.i.d. demonstrations on physical robots, and as a scalable alternative to industry-scale real world evaluations.
@@ 1 Introduction
Simulation is commonly used to evaluate language-guided policies, for example in VLMBench [6], ALFRED [7], CALVIN [8], and VLN [5]. After training a policy on various tasks, the policy is evaluated on a large number of predefined test instances. Since simulations are typically insufficient for truly understanding a policy’s real-world performance [9, 10], and despite correlations between simulation and reality [11], there is a need for an evaluation framework to systematically evaluate language-guided robot policies in the physical world. Since the space of possible robot behaviors in different scenes is large, these approaches must also be efficient with respect to experimenter effort.
@@
We take inspiration from contrast sets [12] and propose perturbation functions to systematically and efficiently evaluate the space of test instances. In particular, we
@@ 2 Background and Related Work
Evaluation of Robot Policies. Simulated environments are often used for evaluating the performance of a real-robot system, such as RoboTHOR [10], CODA [9, 24], and others [25] by recreating a simulated counterpart to a real environment, but show ineffective direct sim2real performance without domain randomization or real-world finetuning strategies. There exist correlations between simulation performance and real-world performance even if they do not exactly match [26, 11]; however there are no guarantees about real-world performance. These works also pre-define a set number of tasks in simulation; but it is not scalable to engineer simulators for every new task. The RoboEval benchmark [27] evaluates the effectiveness of language model-generated programs for robotics by using temporal logic to determine performance. With i.i.d. test sampling, the generalizability of a behavior cloned policy can be bounded [28], but, as the number of possible test instructions and scenes scale, that bounding becomes difficult and time-consuming. We find our work complementary to past work and focus on exploring the space of possible test instructions and scenes in an efficient manner to cover the domain of test instances.
@@ 3 Problem Statement and Notation
The space of discrete language instructions is notoriously large, so these policies are typically evaluated in simulation over thousands of instructions. Several works have have focused on correlating simulation performance to real world performance on a handful of instructions for tuning the sim2real gap [11] or evaluated image-based navigation policies in Airbnbs or rented homes [30, 25]. Tailoring simulations for new skills or renting dozens of environments are not scalable paradigms as experimenters must spend dozens of hours to cover situations policies may encounter.
@@
Evaluation sets. We formalize the problem of evaluating language-guided robot policies in a domain of test instances $\mathcal{X}$ and range of expected behaviors $\mathcal{Y}$. An evaluation set ${X\times Y};{X\subset\mathcal{X}},{Y\subset\mathcal{Y}}$ is composed of instances, the initial conditions $(l,s)\in X$ faced by a policy and the desired outcome $b\in Y$, characterized by the language instruction $l$, the starting scene $s$, and the expected optimal behavior $b$. To evaluate a language-guided robot policy $f$, experimenters sample an evaluation set $X=\{X_{1},...,X_{n}|X_{i}\sim\mathcal{X}\}$ with associated behaviors $Y=\{\text{human}(X_{i})=Y_{i};X_{i}\in X\}$. We assume that the sampled evaluation set $X$ is representative of the test manifold defined by $\mathcal{X}$. A trained, language-guided robot policy $f:\theta\times X\rightarrow Y$ is evaluated over $X$ to produce $\hat{Y}$, against which a performance metric $M(\hat{Y},Y)$ is calculated.
@@
Evaluation Strategies. An evaluation strategy sequentially samples or modifies test instances $X$ into a sequences of test instances to be executed, with experimenter intervention to set up each subsequent starting condition. A standard evaluation strategy $\mathcal{I}(X)=(x_{i}\in X|\forall i\in 1,...,n)$ simply converts the evaluation set into a shuffled sequence of test instances. Let $\mathcal{P}$ be a set of perturbations functions. A strategy $\Gamma$ can construct a larger test set by applying perturbations $\Gamma(X)=(\delta(x)|x_{i}\in X,\delta\in\mathcal{P},\forall i\in{1,...,n})$. In this paper, we highlight the comparative advantages of such a perturbation-based evaluation strategy.
@@
Cost and metrics. To calculate the cost of a series of evaluations, we define an evaluation cost $C(\mathcal{I}(X))$. This cost will give us insight to how much experimenter effort a given evaluation strategy takes. The purpose of a test set is to estimate the value of a given metric $M(f(\mathcal{I}(X)),Y)$ that is an indicator of the robot’s performance. The metric $M$ and evaluation cost $C$ are general terms and is chosen depending on the setting. The metric could be reward in an environment, success rate, success-weighted path length, or the distance from goal, while the evaluation cost metric could be distance of objects moved, or time taken or energy used to reset the scene.
@@
The i.i.d. random sampling strategy $\mathcal{I}({X})$ is typically ideal for estimating an evaluation metric for a given policy. However, the evaluation cost over this strategy $C(\mathcal{I}({X}))$ likely exceeds any given cost budget ${K}$, because i.i.d. sampling of test instances is not aware of the cost of changing instances. The goal of our work is to design a strategy $\Gamma({X})$ over the evaluation set using cost-effective perturbations $\mathcal{P}$ such that $C(\Gamma_{{K}}(X))<C(\mathcal{I}({X}))$ while $M(\Gamma_{{K}}({X}))\approx M(\mathcal{I}({X}))$, where $\Gamma_{{K}}(X)$ is the sequence of evaluations bounded by the cost budget ${K}$.
@@
In this work, we consider the problem of efficiently evaluating language-guided robots across different instructions, scenes and behaviors. As such, we do not focus on object manipulation as the scope of perturbations is limited with pick and place style tasks. We consider settings where it is possible to compose language instructions to define innumerable numbers of robot behaviors.
@@ 4 Contrast Set Evaluation Strategies
Instantiated perturbation functions $\mathcal{P}$ depend on the tasks being evaluated. In our examples, cost $C$ is measured with respect to modifying the environment, so language-based perturbations do not add any additional physical cost. We do not include the cost of resetting the robot position as this is cost is fixed across evaluations, and in some cases can be automated.
@@
Often, when researchers demonstrate the capabilities of their policy, they do not construct an i.i.d. evaluation set, but instead only apply perturbations. These demonstrations are typically done to cheaply evaluate their robot in a limited domain. However, because those test instances and perturbations are not i.i.d., those works are not properly evaluating their robot systems. Since our work assumes access to an i.i.d. test set and then applies perturbations in sequence, contrast sets for robots strikes a balance between saving experimenter effort and properly evaluating a robot policy. Through more systematic evaluations in the test domain, contrast sets also finds better coverage of the test domain, accurate performance estimates, and insights about a policy’s sensitivity to perturbations.
Figure 2: Language-Table Rollouts. In the Language-Table simualtor [31], we sample an evaluation set of 250 test instances that is sequentially evaluated. A test instance is sampled from one of five task types which manipulate blocks according to a task definition. The standard evaluation requires i.i.d. random sampling instructions and scenes, which accumulate more effort for the experimenter. Contrast set evaluation allows experimenters to perturb sampled test instances by making minimal changes after each execution, leading less work for the experimenter.
@@ 5 Language-Table Simulator Experiments
Language-Table [31] is a multi-task language-conditioned control environment and benchmark that spans 696 unique task conditions across five task categories, each of which can be specified in language in dozens of ways. The goal of the tasks is for a manipulator to push blocks of various shapes and colors to different absolute and relative positions on a table, as visualized in Figure 2. It is infeasible to evaluate all of these test instances in a physical environment, especially if there are multiple checkpoints or models to evaluate.
@@ 5.1 Experiment Design
Using this simulator, we test a standard evaluation set ${X}$ against strategies $\Gamma({X})$ that use various perturbation sets $\mathcal{P}$. We sample an evaluation set $X$ from the simulator which consists of 250 test instances equally sampled across five different task categories. Over three seeds, we sample new evaluation sets and run various evaluation strategies.
Perturbations. Using perturbation functions $\mathcal{P}$, we perturb these test instances with low cost. In manipulation, many objects can be interacted with and move around in a scene. We believe it would normally be infeasible to reset the scene after each evaluation, so between perturbed test instances, we do not reset the environment. However, if we applied perturbations that do not modify the expected goal of the robot, a perturbed test instance may already be in a success state. Therefore, we do not use the $\Delta L$ and $\Delta B$ functions. We define two $\Delta LB$ perturbations and two $\Delta SB$ perturbations. As seen in Figure 2, the test instructions typically involve a source block that needs to be moved, a target block that a block needs to move relative to, or a direction such as “top” or “right”. The $\Delta LB_{1}$ perturbation swaps the target and source block referring expressions in the instruction. $\Delta LB_{2}$ modifies the instruction such that any directions or positions are flipped, for example left is changed to right. $\Delta SB_{1}$ moves the target block to a new location, while $\Delta SB_{2}$ moves the source block to a new location. Given a pre-defined test set, we design various contrast sets based on these perturbation strategies to determine their impact on efficiency.
Metrics. In Language-Table, the metric $M(\cdot)$ is the success weighted by path length (SPL). For the cost $C(\cdot)$, we want a measure of the amount of effort an experimenter needs to give in resetting a scene. We define the cost $C(\cdot)$ as the distance in meters of objects moved in the scene.
@@ 5.2 Results and Discussion
Contrast sets can show the policy’s sensitivity todifferent perturbations. Since the perturbations are concrete and specific, they can show a policy’s sensitivity to the language and scene axes. As shown in Figure 3, we find that the language-based perturbation $\Delta LB_{2}$ is notably below the mean SPL for the policy. $\Delta LB_{2}$ simply swaps the directions to be the opposite. If the policy was instructed to move a block to the left of the table, after $\Delta LB_{2}$ was applied, we found the policy would continue to move their block to the right instead. This reduction in performance indicates that the trained model may be overfit to direction in some cases. By contrast, $\Delta SB_{2}$’s higher SPL relative to the full contrast set indicates that the policy is robust to various source block locations. Intuitively, this result shows that certain perturbations help qualitatively characterize the performance of different regions of the test manifold.
Limited interventions are not predictive of SPL. In practice, sometimes experimenters simply execute new language instructions in a given scene until they cannot anymore, which allows them to reduce the amount of effort required by minimally intervening to reset their scene. We explore a limited intervention evaluation strategy in which new language instructions are sampled from the evaluation set sequentially without resetting the scene. However, if a ground truth planner determines an instruction is infeasible, we i.i.d. sample a new scene. As shown in Figure 3, this style of evaluation executes 250 trials with much lower cost than all of the other strategies, but poorly estimates the true test set SPL.
Contrast sets improve predictive performance and provide a cost-effective approach to executing more trials. Using the distance blocks are moved in the scene as a proxy for experimenter effort, Figure 3 shows that language-based perturbations underestimate the SPL of the test set while scene-based perturbations consistently overestimate SPL of the evaluation set. When using all four perturbation functions together, this contrast set has the best predictive performance as it is able to converge closest to the full evaluation set performance. Additionally, the full standard evaluation has a max cost of 281 meters of objects moved; however, contrast sets allows the user to execute more trials for the same cost. Assuming the experimenter has a cost budget of 281 meters of objects moved, standard evaluation allows for 250 trials. The language-based contrast set allows for around 400 trials for the same cost; however, as we have noted, this contrast set under-estimates test performance. The scene and language-based contrast set would contain 380 trials for the same cost without significant under- or over-estimation of the test performance.
@@ 6 VLN-CE Evaluation on a Physical Robot
Vision-and-language navigation in continuous environments [32] (VLN-CE) simulation task where an agent follows fine-grained language instructions in a household setting. There are innumerable ways instructions based on various scenes. We deploy a VLN-CE agent to control a real world Locobot [33]. By evaluating with contrast sets, we are able to draw useful insights about policy sensitivity and estimate the full test performance with a lower cost to the experimenter.

### v2

@@ Abstract
Robot evaluations in language-guided, real world settings are time-consuming and often sample only a small space of potential instructions across complex scenes. In this work, we introduce contrast sets for robotics as an approach to make small, but specific, perturbations to otherwise independent, identically distributed (i.i.d.) test instances. We investigate the relationship between experimenter effort to carry out an evaluation and the resulting estimated test performance as well as the insights that can be drawn from performance on perturbed instances. We use the relative performance change of different contrast set perturbations to characterize policies at reduced experimenter effort in both a simulated manipulation task and a physical robot vision-and-language navigation task. We encourage the use of contrast set evaluations as a more informative alternative to small scale, i.i.d. demonstrations on physical robots, and as a scalable alternative to industry-scale real world evaluations.
@@ 1 Introduction
Simulation is commonly used to evaluate language-guided policies [6, 7, 8, 5]. After training a policy on various tasks, the policy is evaluated on a large number of predefined test instances. Since simulations are typically insufficient for truly understanding a policy’s real-world performance [9, 10], and despite correlations between simulation and reality [11], there is a need for an evaluation framework to systematically evaluate language-guided robot policies in the physical world. Since the space of possible robot behaviors in different scenes is large, these approaches must also be efficient with respect to experimenter effort.
@@
Contrast sets [12] use deliberate perturbations of i.i.d. test sets to enable better evaluation of a machine learning model. We take inspiration from contrast sets and propose perturbation functions to systematically and efficiently evaluate the space of robot test instances. In this work, we build a framework for contrast set-based perturbation strategies for evaluating language-guided robot policies. We apply perturbations on test set instances in the language, scene, and behavior axes, which allows for a lower cost evaluation. This framework strikes a balance between expensive i.i.d. sampling of test instances and cheaper test instance perturbations. In particular, we
@@ 2 Background and Related Work
Evaluation of Robot Policies. Simulation is a common way to train and evaluate robot policies [24, 25, 26]. These simulated environments are often used for evaluating the performance of a real-robot system [10, 9, 27, 28] by recreating a simulated counterpart to a real environment, but show ineffective direct sim2real performance without domain randomization or real-world finetuning strategies. There exist correlations between simulation and real-world performance even if they do not exactly match [29, 11]; however there are no guarantees about real-world performance. These works also pre-define a set number of tasks in simulation; but it is not scalable to engineer simulators for every new task. Other recent work consider how to evaluate in language-guided robots in the real world such as evaluating LLM-based task planners [30] or providing bounds on policy performance [31]. Additionally, RL frameworks such as autonomous RL [32] and others [33, 34] focus on minimizing the number of human interventions during deployment. Contemporary work have begun to investigate how different visual or linguistic attributes impact evaluation performance [35, 36, 37] and efficient data collection [38]. We find our work complementary to these, and focus on evaluating the space of possible test instructions and scenes in an efficient manner.
@@ 3 Problem Statement and Notation
The space of discrete language instructions is notoriously large, so these policies are typically evaluated in simulation over thousands of instructions. Several works have focused on correlating simulation performance to real world performance on a handful of instructions for tuning the sim2real gap [11] or on evaluating image-based navigation policies in Airbnbs or rented homes [40, 28]. Tailoring simulations for new skills or renting dozens of environments are not scalable paradigms as experimenters must spend dozens of hours to cover situations policies may encounter.
@@
Evaluation sets. We formalize the problem of evaluating language-guided robot policies in a domain of test instances $\mathcal{X}$ and range of expected behaviors $\mathcal{Y}$. An evaluation set ${X\times Y};{X\subset\mathcal{X}},{Y\subset\mathcal{Y}}$ is composed of instances, the initial conditions $(l,s)\in X$ faced by a policy and the desired outcome $b\in Y$, characterized by the language instruction $l$, the starting scene $s$, and the expected optimal behavior $b$. For notational simplicity, we specify a language-guided policy $f$ which takes in initial conditions $l,s$ and produces behaviors $b$ rather than that of an iterated state-to-action definition: $f(l,s)\rightarrow b$. To evaluate a language-guided robot policy $f$, experimenters sample an evaluation set $X=\{X_{1},...,X_{n}|X_{i}\sim\mathcal{X}\}$ with associated behaviors $Y=\{\text{human}(X_{i})=Y_{i};X_{i}\in X\}$. We assume that the sampled evaluation set $X$ is representative of the test manifold defined by $\mathcal{X}$. A trained, language-guided robot policy $f:\theta\times X\rightarrow Y$ is evaluated over $X$ to produce $\hat{Y}$, against which a performance metric $M(\hat{Y},Y)$ is calculated.
@@
Evaluation Strategies. An evaluation strategy sequentially samples or modifies test instances $X$ into a sequence of test instances to be executed, with experimenter intervention to set up each subsequent starting condition. A standard evaluation strategy $\mathcal{I}(X)=(x_{i}\in X|\forall i\in 1,...,n)$ simply converts the evaluation set into a shuffled sequence of test instances. Let $\mathcal{P}$ be a set of perturbations functions. A strategy $\Gamma$ can construct a larger test set by applying perturbations $\Gamma(X)=(\delta(x)|x_{i}\in X,\delta\in\mathcal{P},\forall i\in{1,...,n})$. In this paper, we highlight the comparative advantages of such a perturbation-based evaluation strategy.
@@
Cost and metrics. To calculate the cost of a series of evaluations, we define an evaluation cost $C(\mathcal{I}(X))$. This cost will give us insight to how much experimenter effort a given evaluation strategy takes. The purpose of a test set is to estimate the value of a given metric $M(f(\mathcal{I}(X)),Y)$ that is an indicator of the robot’s performance. The metric $M$ and evaluation cost $C$ are general terms and are chosen depending on the setting. The metric could be reward in an environment, success rate, success-weighted path length, or the distance from goal, while the evaluation cost metric could be distance of objects moved, time taken, or energy used to reset the scene.
@@
The i.i.d. random sampling strategy $\mathcal{I}({X})$ is typically ideal for estimating an evaluation metric for a given policy. However, the cost over this strategy $C(\mathcal{I}({X}))$ likely exceeds any given cost budget ${K}$, because i.i.d. sampling of test instances is not aware of the cost of changing instances. The goal of our work is to design a strategy $\Gamma({X})$ over the evaluation set using cost-effective perturbations $\mathcal{P}$ such that $C(\Gamma_{{K}}(X))<C(\mathcal{I}({X}))$ while $M(\Gamma_{{K}}({X}))\approx M(\mathcal{I}({X}))$, where $\Gamma_{{K}}(X)$ is the sequence of evaluations bounded by the cost budget ${K}$. This formulation means that the size of the contrast set is often greater than the size of the standard evaluation $|\Gamma_{{K}}(X)|>|\mathcal{I}({X})|$. The procedure for applying a set of perturbations $\mathcal{P}$ given a cost budget $K$ is shown in Algorithm 1 in the Appendix.
@@
In this work, we consider the problem of efficiently evaluating language-guided robots across different instructions, scenes and behaviors. As such, we do not focus on language-conditioned pick and place tasks as the scope of perturbations is limited in these environments. We consider settings where it is possible to compose language instructions to define innumerable numbers of robot behaviors. Specifically, we use the LanguageTable [41] and VLN-CE [42] tasks, as the language instructions in these environments specifically relate to how the actions are taken.
@@ 4 Contrast Set Evaluation Strategies
Instantiated perturbation functions $\mathcal{P}$ depend on the tasks being evaluated. We measure cost $C$ with respect to modifying the environment, so language-based perturbations do not add any additional physical cost. We do not include the cost of resetting the robot position as it is fixed across evaluations, and in some cases can be automated. Our work investigates whether contrast sets are able to estimate the sample mean performance of a standard evaluation set with a similar or lower cost.
@@
Often, when researchers demonstrate the capabilities of their policy, they do not construct an i.i.d. evaluation set, but instead only apply perturbations. These demonstrations are typically done to cheaply evaluate their robot in a limited domain. However, because those test instances and perturbations are not i.i.d., those works are not properly evaluating their robot systems. Since our work assumes access to an i.i.d. test set and then applies perturbations in sequence, contrast sets for robots strikes a balance between saving experimenter effort and properly evaluating a robot policy. Through more systematic evaluations in the test domain, contrast sets enable better coverage of the test domain, accurate performance estimates, and insights on a policy’s sensitivity to perturbations.
Figure 2: Language-Table Rollouts. In the Language-Table simualtor [41], we sample an evaluation set of 250 test instances that is sequentially evaluated. A test instance is sampled from one of five task types which manipulate blocks according to a task definition. The standard evaluation requires i.i.d. random sampling instructions and scenes, which accumulate more effort for the experimenter. Contrast set evaluation allows experimenters to perturb sampled test instances by making minimal changes after each execution, leading to less work for the experimenter.
@@ 5 Language-Table Simulator Experiments
Language-Table [41] is a multi-task language-conditioned control benchmark that spans 696 unique task conditions across five categories, with each specified in language in dozens of ways. The general task is for a manipulator to push blocks of various shapes and colors to specific relative or absolute positions based on a language instruction, as visualized in Figure 2. It is infeasible to evaluate all of these test instances in a physical environment, especially if there are multiple checkpoints or models to evaluate, making it important to study how to efficiently evaluate large task domains.
@@ 5.1 Experiment Design
We compare a standard evaluation set ${X}$ against strategies $\Gamma({X})$ that use various perturbation sets $\mathcal{P}$. Over three seeds, we sample an evaluation set $X$ from the simulator consisting of 250 test instances sampled across five different task categories, from which we run various evaluation strategies.
Perturbations. Using perturbation functions $\mathcal{P}$, we perturb these test instances with low cost. We believe it would normally be infeasible to reset the scene after each evaluation, so between perturbed test instances, we do not reset the environment. However, if we applied perturbations that do not modify the expected goal of the robot, a perturbed test instance may already be in a success state. Therefore, we do not use the $\Delta L$ and $\Delta B$ functions and define two $\Delta LB$ perturbations and two $\Delta SB$ perturbations as shown in Figure 2. The test instructions typically involve a source block that needs to be moved, a target block that a block needs to move relative to, or a direction such as “top” or “right”. The $\Delta LB_{1}$ perturbation swaps the target and source block referring expressions in the instruction. $\Delta LB_{2}$ modifies the instruction such that any directions or positions are flipped, for example upper left is changed to lower right. $\Delta SB_{1}$ moves the target block to a new location, while $\Delta SB_{2}$ moves the source block to a new location. Given a pre-defined test set, we design various contrast sets from these perturbation strategies to determine their impact on effort.
Metrics. We use an optimal planner to compute the the success weighted by path length (SPL) as the metric $M(\cdot)$ such that suboptimal trajectories are penalized. As a proxy for experimenter effort in resetting a scene, we define the cost $C(\cdot)$ as the distance (in meters) objects are moved in a scene.
@@ 5.2 Results and Discussion
Contrast sets show the policy’s sensitivity to perturbations. Since the perturbations are concrete and specific, they can show a policy’s sensitivity to the language and scene axes. As shown in Figure 3, we find that the language-based perturbation $\Delta LB_{2}$ is notably below the mean SPL for the policy. $\Delta LB_{2}$ simply swaps the directions in text to be the opposite. We took a test instruction of moving a block to the right changed the direction to left with $\Delta LB_{2}$, and we found that the policy would continue to move the block to the right This reduction in performance indicates that the trained model may be overfit to direction in some cases. By contrast, $\Delta SB_{2}$’s higher SPL relative to the contrast set indicates the policy is robust to source block locations. Intuitively, this result shows that perturbations help qualitatively characterize different regions of the test manifold.
Limited interventions are not predictive of SPL. Standard evaluation i.i.d. samples new test instances. However, in practice, sometimes experimenters simply execute new language instructions in a given scene until they cannot anymore, which allows them to reduce the amount of effort required by minimally intervening to reset their scene. We explore a limited intervention evaluation strategy in which new language instructions are sampled from the evaluation set sequentially without resetting the scene. This experiment is meant to be a lower bound in cost, where new language instructions and behaviors are sampled, but the reset costs are is minimized by not resetting the scene unless a planner determines an instruction is infeasible. As shown in Figure 3, 250 trials of this strategy yields much lower cost than all of the other strategies, but poorly estimates the test SPL.
Contrast sets improve predictive performance and provide a cost-effective approach to executing more trials. Figure 3 shows that language-based perturbations underestimate the SPL of the evaluation set while scene-based perturbations overestimates it. Using all four perturbation functions together yields the best predictive performance, as it converges closest to the full evaluation set performance. As shown in Figure 6 in the Appendix, the standard evaluation incurs a cost of 281 meters over 250 trials. In comparison, the language-based contrast set reaches around 400 trials for the same cost but under-estimates test performance. The combined scene and language-based contrast set allows for 380 trials without significant under- or over-estimation of the test performance.
@@ 6 VLN-CE Evaluation on a Physical Robot
Vision-and-language navigation in continuous environments [42] (VLN-CE) is a task where an agent follows fine-grained language instructions in a household setting. We deploy a VLN-CE agent to control a real world Locobot [43]. By evaluating with contrast sets, we are able to draw insights about policy sensitivity and estimate the full test performance with a lower cost to the experimenter.